Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

ODDS

Expert replies
by scud » Sun Oct 05, 2008 2:22 am
Horse Name
Odds of Placing in the top three

The Baron 3/5


Happy Cynic 7/10


California Girl 3/4


Love Song 1/4


Inamorata 1/2



The chart above lists the odds that a horse will place in the top three. As part of a contest, if at least 2 of the horses that a person bets upon finish in the top three, the bettor receives a T-shirt. If Cecilia bets on The Baron, Happy Cynic, and Inamorata, what is the probability that she does not receive a T-shirt?


A) 3/50
B)21/100
C)29/100
D)35/100
E)65/100
Join the discussion
Source: — Problem Solving |

by Morgoth » Sun Oct 05, 2008 3:07 am
probability that cecilia does not receive t-shirt

= only 1 horse finish in top 3

only 1 horse finish in top 3 [we can have 3 cases]

Barron wins, other two lose
3/5*3/10*1/2 = 9/100

Happy Cynic wins, other two lose
7/10*1/2*2/5 = 7/50

Inamorata wins, other two loses
3/10*2/5*1/2 = 3/50

Combine cases = 9/100 + 7/50 + 3/50 = (9+14+6)/100 = 29/100.

OA?
Last edited by Morgoth on Sun Oct 05, 2008 8:12 am, edited 1 time in total.
Join the discussion

by scud » Sun Oct 05, 2008 7:31 am
first of all thanks for replying....but it seems you have not read the question prpoperly, it says that only 5 horses are there and 3 will win so the case of all three horses loosing can not be possible......but yeah can you explain me the meaning of 'odds of placing in top three'....i am confused what it means prob of winning or of loosing
Join the discussion

by Stockmoose16 » Mon Oct 06, 2008 5:15 pm
Morgoth wrote:probability that cecilia does not receive t-shirt

= only 1 horse finish in top 3

only 1 horse finish in top 3 [we can have 3 cases]

Barron wins, other two lose
3/5*3/10*1/2 = 9/100

Happy Cynic wins, other two lose
7/10*1/2*2/5 = 7/50

Inamorata wins, other two loses
3/10*2/5*1/2 = 3/50

Combine cases = 9/100 + 7/50 + 3/50 = (9+14+6)/100 = 29/100.

OA?
I believe this is correct, except the question stem says the ODDS of a horse placing in the top 3, while the actual question asks for the PROBABILITY.

I think the question is worded poorly, because, for example in poker, if your odds of winning a hand are 2/1, that means the probability is 1/3. Can someone confirm my logic here.
Join the discussion

by Stockmoose16 » Mon Oct 06, 2008 5:30 pm
Morgoth wrote:probability that cecilia does not receive t-shirt

= only 1 horse finish in top 3

only 1 horse finish in top 3 [we can have 3 cases]

Barron wins, other two lose
3/5*3/10*1/2 = 9/100

Happy Cynic wins, other two lose
7/10*1/2*2/5 = 7/50

Inamorata wins, other two loses
3/10*2/5*1/2 = 3/50

Combine cases = 9/100 + 7/50 + 3/50 = (9+14+6)/100 = 29/100.

OA?
This question is written incorrectly, I believe. To answer the question above, you'd need to change ODDS into PROBABILITY (two different things). So, for example, the odds of Barron coming in the top 3 are 3:5, however, the probability of Barron coming in the top 3 is 3/8 (you must add the outcomes together).

Therefore the correct answer to this question should be the following:

(1) Barron in top 3, other 2 picks not: 3/8 * 10/17 *2/3= 60/408
+
(2) HC in top 3, other 2 picks not: 5/8 * 7/17 * 2/3 =70/408
+
(3) Imarota in top 3, other 2 picks not: 5/8*10/17*1/3=50/408

answer is: 180/408= 90/204= 45/102

I'm sure an expert will confirm that my answer is correct, and the problem is worded incorrectly... someone please comment.
Join the discussion